Find Vf at the given point. f(x,y,z)=e*** cos z + (y + 2) sinx (Type an exact answer, using radicals as needed.)
Answer #1

(3,1,2)= ext=cos Z +(4+2) Sinter) Now, he e ex++ cos2 + (3+2) firme So o 1 (0.0, I = COS I +2 = tz +2 = 472 of extCsZ + Sint